The financial statements for Castile Products, Inc., are given below:

Castile Products, Inc.
Balance Sheet
December 31
Assets
Current assets:
Cash $ 21,000
Accounts receivable, net 170,000
Merchandise inventory 390,000
Prepaid expenses 9,000
Total current assets 590,000
Property and equipment, net 840,000
Total assets $ 1,430,000
Liabilities and Stockholders' Equity
Liabilities:
Current liabilities $ 230,000
Bonds payable, 12% 300,000
Total liabilities 530,000
Stockholders’ equity:
Common stock, $5 par value $ 160,000
Retained earnings 740,000
Total stockholders’ equity 900,000
Total liabilities and stockholders’ equity $ 1,430,000
Castile Products, Inc.
Income Statement
For the Year Ended December 31
Sales $ 2,300,000
Cost of goods sold 1,170,000
Gross margin 1,130,000
Selling and administrative expenses 630,000
Net operating income 500,000
Interest expense 36,000
Net income before taxes 464,000
Income taxes (30%) 139,200
Net income $ 324,800

Account balances at the beginning of the year were: accounts receivable, $210,000; and inventory, $300,000. All sales were on account. Assets at the beginning of the year totaled $1,000,000, and the stockholders’ equity totaled $685,000.

Required:

Compute the following: (For Requirements 1 to 4, enter your percentage answers rounded to 2 decimal places (i.e., 0.1234 should be entered as 12.34).)

1. Gross margin percentage.

2. Net profit margin percentage.

3. Return on total assets.

4. Return on equity.

5. Was financial leverage positive or negative for the year?
Answer #1

Ans:

1) Gross margin percentage = Gross profit/Sales =1,130,000/2,300,000= 49.13%
2) Net profit margin percentage = Net income/Sales = 324,800/2,300,000= 14.13%
3) Return on total assets = Net Income / AVERAGE TOTAL ASSETS
AVERAGE TOTAL ASSETS = ( $1,000,000+ $1,430,000 ) / 2 = 1215000
324800 / 1215000 = 26.73%

4) Return on equity = Net Income/Average share holders equity=324800/792500=40.98%
AVERAGE TOTAL EQUITY =
900000+685,000=792,500

5) Financial leverage = Average total assets/average equity
=1215000/792500=1.533